# MoGenTS: Motion Generation based on Spatial-Temporal Joint Modeling

### v1

@@ 2.1 Motion Generation from Continuous Regression
The initial methods devised for human motion generation were predominantly focused on directly regressing continuous motion values, a rather straightforward approach without intermediary quantization [2, 3, 25, 26, 27, 28]. Numerous strategies have utilized the variational autoencoder (VAE) framework to merge the latent embeddings of encoded text with those of encoded poses, subsequently decoding this combined representation into motion predictions [2, 3, 26, 4, 5]. Others have explored the potential of the recurrent network [25, 29, 30], generative adversarial network [28, 31, 1], or transformer network [32, 33, 34, 35] to improve the quality of motion regression. Drawing on the success of diffusion models, recent methods have begun to introduce the diffusion process into motion diffusion [36, 6, 7, 8, 9, 10, 11, 37, 38], yielding commendable results. To improve the diffusion model in human motion generation, various mechanisms are proposed to be integrated into the denoising process. A hybrid retrieval mechanism considering both semantic and kinematic similarities is integrated to refine the denoising process in [10]. Physical constraints [37] and guidance constraints [39] are also incorporated into the diffusion process for more reasonable motions. In addition, a module of construction supported by linguistics structure that constructs accurate and complete language characteristics and a module of progressive reasoning aware of the context are designed in [38].
@@ 3.1 Problem Statement
Motion Quantization. Given a motion sequence $\{\mathbf{M}_{t}\}_{t=1}^{T}$, the objective is to construct an optimal codebook $\mathbf{C}=\{\mathbf{c}_{i}\}_{i=1}^{C}$ capable of accurately representing this sequence, so that the encoded motion features $\mathbf{V}$ could be substituted by the corresponding codes $c_{i}$ from the codebook with the minimum loss, resulting in a discrete token sequence $\{\mathbb{O}\}$, representing the input motion.
Motion Generation. Given the textual prompt $\mathbf{T}$, the task of our framework is to generate the indices of the tokens $\{\mathbb{O}\}$, so that the vectors corresponding to these tokens could be decoded into a motion sequence aligned with the input text.
@@ 3.2 Overview
The overview of our method is illustrated in Figure 1. Our framework could be divided into two parts, motion quantization and motion generation. In the quantization phase, we first represent the input motion sequence in the joint-time structure, and encode it into a 2D vector map. Subsequently, the vectors are quantized into codes of a codebook, represented by the indices of the selected codebook entries, i.e., the joint tokens. Therefore, after the quantization, the input motion sequence is converted to tokens arranged in a 2D structure, where one dimension is spatial while the other one is temporal. This results in a 2D motion map, which is similar to a 2D image. In the generation phase, we mask the 2D token map with a temporal-spatial 2D masking strategy, and then use a 2D transformer to predict the masked tokens, conditioned on the embedding $\mathbf{T}$ of the given text prompt. The 2D motion transformer considers both spatial attention and temporal attention between different 2D tokens. The 2D position embedding $\mathbf{P}$ is also used to convey the spatial and temporal locations of each token. Finally, the generated tokens are decoded back into a motion sequence that aligns with the given textual prompt, thus completing the process of text-driven motion generation.
@@ 3.3 Spatial-temporal 2D Joint Quantization of Motion
Directly regressing the continuous motion sequence is not easy. Thus, we perform motion quantization following previous methods [17] to convert the regression problem to a classification problem, but differently, we quantize each joint rather than the whole-body pose. Therefore, we organize the motion sequence in a 2D structure and process it as a 2D map, as illustrated in Figure 2. For a motion with sequence length $T$ and joint number $J$, we employ a 2D convolutional network to encode the motion joints $\mathbf{J}=\{\mathbf{j}_{t}^{j}\}_{t=1,...,T}^{j=1,...,J}$ into 2D vectors $\mathbf{V}=\{\mathbf{v}_{t}^{j}\}_{t=1,...,T}^{j=1,...,J}$, where each vector $\mathbf{v}_{t}^{j}$ corresponds to time $t$ and joint $j$, as
@@
where $\mathcal{E}$ denotes the encoder network composed of two convolutional residual blocks. Then we quantize the vector with a codebook, i.e. we replace the vector $\mathbf{v}_{t}^{j}$ with its nearest code entry $\mathbf{\tilde{v}}_{t}^{j}$ in a preset codebook $\mathbf{C}=\{\mathbf{c}_{i}\}_{i=1}^{C}$, as
@@
where $\alpha$ denotes a weighting factor. We also employ the residual VQ-VAE structure, exponential moving average, and codebook reset following previous methods [14, 17].
@@ 3.5.1 Spatial-Temporal Motion Transformer
Spatial-Temporal 2D Attention. For the spatial-temporal attention, we first extract the text feature embedding using CLIP [48], resulting in a text token $\mathbb{O}_{\text{text}}$. Then we add a 2D position encoding [49] to the 2D token map. The position encoding $\mathbf{P}_{t}^{j}$ is computed by the sinusoidal function in the spatial dimension $j$ and in the temporal dimension $t$, respectively, so it provides both the spatial position and the temporal position information for the subsequent attention network. After the 2D position embedding, we flatten the 2D token map to the 1D structure and prepend the text embedding to the flattened token sequence. Then we perform the 1D attention in the flattened spatial-temporal dimension, where the spatial and temporal relationship is reserved by the 2D position encoding, as
@@
Joint Spatial Attention. Joint spatial attention is the one-dimensional attention in only the spatial dimension. As shown in Figure 3(b), we ignore the temporal dimension and rearrange the 2D tokens. The tokens of different time frames are regarded as different batches, and then the attention is only calculated between different joints on each time batch, as
@@
Joint Temporal Attention. Similar to spatial attention, we also perform one-dimensional attention in only the temporal dimension. As shown in Figure 3(b), we ignore the spatial dimension and rearrange the 2D tokens. The tokens of different joints are regarded as different batches, and then the attention is only calculated between different times on each joint batch, as
@@
Integration of attention modules. Our approach applies spatial-temporal 2D attention, joint spatial attention, and joint temporal attention in a sequential manner. Here the spatial-temporal 2D attention provides the initialization, and joint spatial attention and joint-temporal attention offer regualarization.
@@
Training. In the training, we use the spatial-temporal motion transformer to predict the probability logits of all $C$ tokens $\hat{y}_{i}$ in the masked positions. Then a cross-entropy loss is computed between the prediction $\hat{y}$ and the target $y$, as
@@
Inference. In the inference, the generation starts from an empty token map, i.e., all tokens are masked. The prediction of the masked tokens is repeated by $N$ iterations. In each iteration, we build the probability distribution of $C$ tokens from the SoftMax of the predicted logits, and sample the output from the distribution. In the next iteration, according to the schedule shown in Section 3.4 with $\tau=n/N$, low-score tokens are selected to be masked and predicted again, until $n$ reaches $N$. Once the iterative masked generation is finished, the residual transformer progressively predicts the residual tokens of the residual VQ layers. Afterward, the base tokens plus the residual tokens are decoded by the VQVAE decoder to get the final generated motion.
@@ 4.2 Implementation details
Our framework is trained on an NVIDIA A100 GPU with PyTorch. The batch size is set to 256 and the learning rate is set to 2e-4. To quantize the motion data into our 2D structure, we restructure the pose in the datasets to a joint-based format, with the size of $12\times J$. The data is then represented by the joint VQ codebook comprised of 256 codes, each with a dimension of 1024. Due to the global information included in the HumanML3D data representation, we also incorporate an additional global VQ codebook of the same size to encode the global information. For the residual structure, the number of residual layers is set to 5 following previous methods [17]. Both the encoder and decoder are constructed from 2 convolutional residual blocks with a downscale of 4. The transformers in our model are all set to have 6 layers, 6 heads, and 384 latent dimensions. The parameter $\alpha$ is set to 1 and $N$ is set to 10. More details are described in the supplementary material.
@@ 4.3 Evaluation
Evaluation of Motion Quantization. The core idea of our framework starts from the spatial-temporal 2D joint quantization. Therefore, we first evaluate our VQVAE against that of previous quantization-based methods. For a fair comparison, we temporarily configure our codebook to the same size as the previous methods [17], i.e., $512\times 512$. As reported in Table 2, the FID and MPJPE accuracy of our method exceeds previous methods by a large margin, in both the HumanML3D and the KIT-ML datasets. This reveals that the difficulty of quantizing an individual joint is much smaller than quantizing a whole-body pose.
Evaluation of Motion Generation. We then compare our model to previous text-to-motion works, including both the continuous regression-based methods and the discrete quantization-based methods. From the results reported in Table 1, our method outperforms all previous methods on both the HumanML3D and the KIT-ML datasets, which demonstrates the effectiveness of our method. Specifically, the FID is decreased by $26.6\%$ on HumanML3D and is decreased by $29.9\%$ on KIT-ML. In addition, the R-precision even significantly surpasses the ground truth. From the qualitative results displayed in Figure 4, we also see that our method generates motions that are better matched with the input texts compared to previous methods.
@@ 4.4 Ablation Study
Joint 2D VQ. To verify the efficacy of the proposed framework, we first establish a baseline model incorporating a 1D motion VQVAE and a 1D token prediction transformer. Then we change the VQVAE to our 2D joint VQVAE, examining its impact on both motion auto-encoding and motion generation capabilities. According to the outcomes presented in Table 2, we observed a substantial enhancement in auto-encoding performance. However, as indicated in Table 3, the motion generation performance deteriorates. This decline can be attributed to the considerable increase in the number of tokens, which renders the 1D masking strategy and 1D transformer no longer adequate.
2D Masking. We then change the 1D masking to temporal-spatial 2D token masking, which leads to a significant performance improvement, as depicted in Table 3. This demonstrates the necessity of our 2D masking strategy, as discussed in Section 3.4.
2D Position Encoding. With the 2D VQ, the attention is calculated between 2D tokens, but it is not a full version of the 2D attention. We upgrade this to spatial-temporal 2D attention by integrating the 2D position encoding, which provides both the spatial location and temporal location information. The performance improvement brought by this encoding is recorded in Table 3.
Joint Spatial and Temporal Attention. As discussed in Section 3.5.1, the mix of spatial and temporal information is complex for the bidimensional attention performed on a long sequence. Here we inspect the benefits brought by the individual joint spatial and temporal attention mechanism. From the results in Table 3, this design further improves the performance of our framework.
@@ 5 Conclusion
Limitations and future work. Despite the improvement in the quantization, there is still the approximation error. A large dataset for pretraining an accurate quantizer is needed in future work.
@@ A.1 Evaluation Metrics
where $\mu_{\text{gen}}$ and $\mu_{\text{gt}}$ are the mean of $\mathbf{f}_{\text{gen}}$ and $\mathbf{f}_{\text{gt}}$, $Cov$ denotes the covariance matrix, and $Tr$ denotes the trace of a matrix.
@@
A.2 More Details
Classifier Free Guidance.
where $s$ is set to 4.

### v2

@@ 2.1 Motion Generation from Continuous Regression
The initial methods devised for human motion generation were predominantly focused on directly regressing continuous motion values, a rather straightforward approach without intermediary quantization [2, 3, 25, 26, 27, 28]. Numerous strategies have utilized the variational autoencoder (VAE) framework to merge the latent embeddings of encoded text with those of encoded poses, subsequently decoding this combined representation into motion predictions [2, 3, 26, 4, 5]. Others have explored the potential of the recurrent network [25, 29, 30], generative adversarial network [28, 31, 1], or transformer network [32, 33, 34, 35] to improve the quality of motion regression. Drawing on the success of diffusion models, recent methods have begun to introduce the diffusion process into motion diffusion [36, 6, 7, 8, 9, 10, 11, 37, 38, 39, 40], yielding commendable results. To improve the diffusion model in human motion generation, various mechanisms are proposed to be integrated into the denoising process. A hybrid retrieval mechanism considering both semantic and kinematic similarities is integrated to refine the denoising process in [10]. Physical constraints [37] and guidance constraints [41] are also incorporated into the diffusion process for more reasonable motions. In addition, a module of construction supported by linguistics structure that constructs accurate and complete language characteristics and a module of progressive reasoning aware of the context are designed in [38].
@@ 3.1 Problem Statement
Motion Quantization.
Given a motion sequence $\{\mathbf{M}_{t}\}_{t=1}^{T}$, the objective is to construct an optimal codebook $\mathbf{C}=\{\mathbf{c}_{i}\}_{i=1}^{C}$ capable of accurately representing this sequence, so that the encoded motion features $\mathbf{V}$ could be substituted by the corresponding codes $c_{i}$ from the codebook with the minimum loss, resulting in a discrete token sequence $\{\mathbb{O}\}$, representing the input motion.
Motion Generation.
Given the textual prompt $\mathbf{T}$, the task of our framework is to generate the indices of the tokens $\{\mathbb{O}\}$, so that the vectors corresponding to these tokens could be decoded into a motion sequence aligned with the input text.
@@ 3.2 Overview
The overview of our method is illustrated in Figure 1. Our framework could be divided into two parts, motion quantization and motion generation. In the quantization phase, we first represent the input motion sequence in the joint-time structure, and encode it into a 2D vector map. Subsequently, the vectors are quantized into codes of a codebook, represented by the indices of the selected codebook entries, i.e., the joint tokens. Therefore, after the quantization, the input motion sequence is converted to tokens arranged in a 2D structure, where one dimension is spatial while the other one is temporal. This results in a 2D motion map, which is similar to a 2D image. In the generation phase, we mask the 2D token map with a temporal-spatial 2D masking strategy, and then use a 2D transformer to predict the masked tokens, conditioned on the CLIP [46] embedding $\mathbf{T}$ of the given text prompt. The 2D motion transformer considers both spatial attention and temporal attention between different 2D tokens. The 2D position embedding $\mathbf{P}$ is also used to convey the spatial and temporal locations of each token. Finally, the generated tokens are decoded back into a motion sequence that aligns with the given textual prompt, thus completing the process of text-driven motion generation.
@@ 3.3 Spatial-temporal 2D Joint Quantization of Motion
Directly regressing the continuous motion sequence is not easy. Thus, we perform motion quantization following previous methods on the whole pose [5, 17] or body part [42, 47, 48] to convert the regression problem to a classification problem, but differently, we quantize each joint rather than the whole-body pose. Therefore, we organize the motion sequence in a 2D structure and process it as a 2D map, as illustrated in Figure 2. For a motion with sequence length $T$ and joint number $J$, we employ a 2D convolutional network to encode the motion joints $\mathbf{J}=\{\mathbf{j}_{t}^{j}\}_{t=1,...,T}^{j=1,...,J}$ into 2D vectors $\mathbf{V}=\{\mathbf{v}_{t}^{j}\}_{t=1,...,T}^{j=1,...,J}$, where each vector $\mathbf{v}_{t}^{j}$ corresponds to time $t$ and joint $j$, as
@@
where $\mathcal{E}$ denotes the encoder network composed of two convolutional residual blocks, and $\mathbf{j}_{t}^{j}$ denotes a 3D rotation angle of joint $j$ at time $t$. Then we quantize the vector with a codebook, i.e. we replace the vector $\mathbf{v}_{t}^{j}$ with its nearest code entry $\mathbf{\tilde{v}}_{t}^{j}$ in a preset codebook $\mathbf{C}=\{\mathbf{c}_{i}\}_{i=1}^{C}$, as
@@
where $\alpha$ denotes a weighting factor. We also employ the residual VQ-VAE structure, exponential moving average, and codebook reset following previous methods [14, 17], but for simplicity we only describe our method in single-level quantization without incorporating a residual structure in this section.
@@ 3.5.1 Spatial-Temporal Motion Transformer
Spatial-Temporal 2D Attention.
For the spatial-temporal attention, we first extract the text feature embedding using CLIP [46], resulting in a text token $\mathbb{O}_{\text{text}}$. Then we add a 2D position encoding [53] to the 2D token map. The position encoding $\mathbf{P}_{t}^{j}$ is computed by the sinusoidal function in the spatial dimension $j$ and in the temporal dimension $t$, respectively, so it provides both the spatial position and the temporal position information for the subsequent attention network. After the 2D position embedding, we flatten the 2D token map to the 1D structure. Then we perform the 1D attention in the flattened spatial-temporal dimension, where the spatial and temporal relationship is reserved by the 2D position encoding, as
@@
Joint Spatial Attention.
Joint spatial attention is the one-dimensional attention in only the spatial dimension. As shown in Figure 3(b), we ignore the temporal dimension and rearrange the 2D tokens. The tokens of different time frames are regarded as different batches, and then the attention is only calculated between different joints on each time batch, as
@@
Joint Temporal Attention.
Similar to spatial attention, we also perform one-dimensional attention in only the temporal dimension. As shown in Figure 3(b), we ignore the spatial dimension and rearrange the 2D tokens. The tokens of different joints are regarded as different batches, and then the attention is only calculated between different times on each joint batch, as
@@
Integration of attention modules.
Our approach applies spatial-temporal 2D attention, joint spatial attention, and joint temporal attention in a sequential manner. Here the spatial-temporal 2D attention provides the initialization, and joint spatial attention and joint-temporal attention offer regualarization.
@@
Training.
In the training, we use the spatial-temporal motion transformer to predict the probability logits of all $C$ tokens $\hat{y}_{i}$ in the masked positions. Then a cross-entropy loss is computed between the prediction $\hat{y}$ and the target $y$, as
@@
Inference.
In the inference, the generation starts from an empty token map, i.e., all tokens are masked. The prediction of the masked tokens is repeated by $N$ iterations. In each iteration, we build the probability distribution of $C$ tokens from the SoftMax of the predicted logits, and sample the output from the distribution. In the next iteration, according to the schedule shown in Section 3.4 with $\tau=n/N$, low-score tokens are selected to be masked and predicted again, until $n$ reaches $N$. Once the iterative masked generation is finished, the residual transformer progressively predicts the residual tokens of the residual VQ layers. Afterward, the base tokens plus the residual tokens are decoded by the VQVAE decoder to get the final generated motion.
Classifier Free Guidance.
where $s$ is set to 4.
@@ 4.2 Implementation details
Our framework is trained on two NVIDIA A100 GPUs with PyTorch. The batch size is set to 256 and the learning rate is set to 2e-4. To quantize the motion data into our 2D structure, we restructure the pose in the datasets to a joint-based format, with the size of $12\times J$. The data is then represented by the joint VQ codebook comprised of 256 codes, each with a dimension of 1024. Due to the global information included in the HumanML3D data representation, we also incorporate an additional global VQ codebook of the same size to encode the global information. For the residual structure, the number of residual layers is set to 5 following previous methods [17]. Both the encoder and decoder are constructed from 2 convolutional residual blocks with a downscale of 4. The transformers in our model are all set to have 6 layers, 6 heads, and 384 latent dimensions. The parameter $\alpha$ is set to 1 and $N$ is set to 10. More details are described in the supplementary material.
@@ 4.3 Evaluation
Evaluation of Motion Quantization.
The core idea of our framework starts from the spatial-temporal 2D joint quantization. Therefore, we first evaluate our VQVAE against that of previous quantization-based methods. For a fair comparison, we temporarily configure our codebook to the same size as the previous methods [17], i.e., $512\times 512$. As reported in Table 2, the FID and MPJPE accuracy of our method exceeds previous methods by a large margin, in both the HumanML3D and the KIT-ML datasets. This reveals that the difficulty of quantizing an individual joint is much smaller than quantizing a whole-body pose.
Evaluation of Motion Generation.
We then compare our model to previous text-to-motion works, including both the continuous regression-based methods and the discrete quantization-based methods. From the results reported in Table 1, our method outperforms all previous methods on both the HumanML3D and the KIT-ML datasets, which demonstrates the effectiveness of our method. Specifically, the FID is decreased by $26.6\%$ on HumanML3D and is decreased by $29.9\%$ on KIT-ML. In addition, the R-precision even significantly surpasses the ground truth. From the qualitative results displayed in Figure 4, we also see that our method generates motions that are better matched with the input texts compared to previous methods.
@@ 4.4 Ablation Study
Joint 2D VQ.
To verify the efficacy of the proposed framework, we first establish a baseline model incorporating a 1D motion VQVAE and a 1D token prediction transformer. Then we change the VQVAE to our 2D joint VQVAE, examining its impact on both motion auto-encoding and motion generation capabilities. According to the outcomes presented in Table 2, we observed a substantial enhancement in auto-encoding performance. However, as indicated in Table 3, the motion generation performance deteriorates. This decline can be attributed to the considerable increase in the number of tokens, which renders the 1D masking strategy and 1D transformer no longer adequate.
2D Masking.
We then change the 1D masking to temporal-spatial 2D token masking, which leads to a significant performance improvement, as depicted in Table 3. This demonstrates the necessity of our 2D masking strategy, as discussed in Section 3.4.
2D Position Encoding.
With the 2D VQ, the attention is calculated between 2D tokens, but it is not a full version of the 2D attention. We upgrade this to spatial-temporal 2D attention by integrating the 2D position encoding, which provides both the spatial location and temporal location information. The performance improvement brought by this encoding is recorded in Table 3.
Joint Spatial and Temporal Attention.
As discussed in Section 3.5.1, the mix of spatial and temporal information is complex for the bidimensional attention performed on a long sequence. Here we inspect the benefits brought by the individual joint spatial and temporal attention mechanism. From the results in Table 3, this design further improves the performance of our framework.
@@ 5 Conclusion
5.1 Limitations and future work.
Quantization.
Despite the improvement of our method in the quantization, there is still the approximation error, which limits the motion generation. This should be further improved in the future network design, e.g., employing the coarse-to-fine technology. Another important point is the restricted size of the current human motion dataset. In the image pretraining, a much larger dataset is utilized to achieve decent quantization results. Similarly, a large dataset for pretraining an accurate human motion quantizer is also needed in future work.
Generation.
We follow the mask-and-generation for motion generation in this work. The masking strategy used in the current framework is borrowed from Bert [24]. On the contrary, some other methods follow the auto-regressive generation and also achieve good performance. In the future, one promising direction is combining these two technologies.
@@ A.1 Evaluation Metrics
where $\mu_{\text{gen}}$ and $\mu_{\text{gt}}$ are the mean of $\mathbf{f}_{\text{gen}}$ and $\mathbf{f}_{\text{gt}}$, $Cov$ denotes the covariance matrix, and $Tr$ denotes the trace of a matrix. The features are extracted using the models following previous methods [5].
@@
A.2 Computational Overhead
We test the computational overhead of different methods on an NVIDIA 4090 GPU and report the average inference time per sentence in Table 4. Although our method does not achieve the shortest inference time, the increase in computational overhead is not significant. Overall, the computational overhead of our method is comparable to that of mainstream methods.
Method Average Inference Time per Sentence
MLD 134 ms
MotionDiffuse 6327 ms
MDM 10416 ms
T2M-GPT 239 ms
MoMask 73 ms
Ours 181 ms
Table 4: Computational overhead of different methods.